OpenStudy (anonymous):

f(x)=x^2+x and g(x)=2/(x+3) find value from : a. f(-6) b. g(1/2)

OpenStudy (akshay_budhkar):

a.30

OpenStudy (akshay_budhkar):

b.4/7....

OpenStudy (anonymous):

about the second 4/7.... can u describe...?

OpenStudy (akshay_budhkar):

yes sure 2/(1/2+3)= 2/(7/2)=4/7.. :)

OpenStudy (anonymous):

a. f(-6) f(x)=x^2+x = (-6)^2+(-6) =36-6 =30

OpenStudy (anonymous):

b. g(1/2) g(x)=2/(1/2)+3) = 2 / 7/2 =2*(2/7) = 4/7